OpenStudy (anonymous):

HELP ! :/ Solve the solution using Substitution. Check Your Answer. y = -2x + 8 5y - x + 4 = 0 (Answer must be ordered pair)

OpenStudy (freckles):

have you tried putting in the equation solved for y into the other equation?

OpenStudy (freckles):

like so: 5(-2x+8)-x+4=0 now solve for x

OpenStudy (anonymous):

Ohhhh okay! Give me a moment....

OpenStudy (anonymous):

x = 34? @freckles

OpenStudy (freckles):

so you distribute: -10x+40-x+4=0 combined like terms -11x+44=0 <--- hmmm so looking at this I think you should try again

OpenStudy (anonymous):

So, I'm still looking for X correct?

OpenStudy (freckles):

yep

OpenStudy (freckles):

you might want to look at the equation you "solved" to get x=34 because I do not believe that is correct. I actually showed you to distribute and then combine like terms above. You are actually one step away from solving for x since I have done most of the work for you.

OpenStudy (freckles):

-11x+44=0 so x=?

OpenStudy (anonymous):

4??

OpenStudy (freckles):

sounds good

OpenStudy (freckles):

now do you know how to find y?

OpenStudy (freckles):

y = -2x + 8 you have x is 4 so replace x with 4 this gives you enough info to find y

OpenStudy (anonymous):

So -2x(4) + 8?

OpenStudy (freckles):

yep

OpenStudy (freckles):

where x means multiply right?

OpenStudy (anonymous):

Right @freckles